A camera lens used for taking close-up photographs has a focal length of 20.5 mm. The farthest it can be placed from the film is 31.0 mm.

(a) What is the closest object that can be photographed?

(b) What is the magnification of this closest object?
Answer #1

a) 1/f = 1/Do + 1/Di

1/20.5 = 1/31 + 1/Di

Di = 60.5 mm

b) Magnification = - Di/Do = - 60.5/31 = -1.95

Negative sign indicates inverted image